OpenStudy (anonymous):

give the domain of f(x)=(x+1)^5+1 and the domain of f^-1(x)

OpenStudy (anonymous):

domain of f(x) is all real numbers domain of f(x)^-1 is all real numbers since the power is odd.

OpenStudy (anonymous):

first is all real numbers, second is all real numbers not equal to 0, because it takes values of x and turns them into 1/0 which is undefined

OpenStudy (anonymous):

it turns x into 1/x, which in the case of 0 would be 1/0, I should have said

OpenStudy (anonymous):

domain for the second one is all real nos except -2

OpenStudy (anonymous):

\[f ^{-1}=(\sqrt[5]{x-1})-1\] Domain is all reall numbers since the fifth root exist for all real numbers

OpenStudy (anonymous):

i think hes asking for the reciprocal of f(x) and not the inverse

OpenStudy (anonymous):

i believe xon is correct, its all real numbers for both f and f^-1, you can take 5th roots of any number, positive or negative

OpenStudy (anonymous):

so what is f^-1(0)

OpenStudy (anonymous):

using xon's equation: \[f^{-1}(0) = \sqrt[5]{(0-1) } - 1 = -1-1 = -2\]

OpenStudy (anonymous):

f^-1(0)=(-1)^(1/5)-1 f^-1(0)=-1-1 f^-1(0)=-2

OpenStudy (anonymous):

I assumed it was two seperate unrelated equations, asking for domain of each

OpenStudy (anonymous):

ahh, now I see, no = sign in second

OpenStudy (anonymous):

i guess it really depends on what the poster means by f^-1

OpenStudy (anonymous):

it is the inverse

OpenStudy (anonymous):

cool, that what I thought. go with xon's answer.

OpenStudy (anonymous):

yep, sorry for the confusion

OpenStudy (anonymous):

its all good, nothing like some mathematical debating to shake things up :)

OpenStudy (anonymous):

ya, if you wanted the reciprocal, then it should be 1/f(x) or f(x)^-1 because the inverse is defined as f^-1(x). 1/f(x) is not always equal to f^-1(x)